OpenStudy (anonymous):

Find the roots. 2x^2 + 4x - 5 =0

OpenStudy (anonymous):

2x^2+4x-5=0 x^2+2x-5/2=0 x^2+2x+1-1-5/2=0 (x+1)^2=1+5/2=7/2 x+1=\[\pm \sqrt{\frac{ 7 }{ 2 }}\] x=-1\[\pm \frac{ 7 }{ 2 }\]